OpenStudy (anonymous):

sec540 degrees without using a caculator. Help!!!

OpenStudy (anonymous):

540-360=180 sec = 1/cos cos(180) = -1 sec(180) = 1/cos(180) = 1/-1 = -1

OpenStudy (abdullah1995):

pls dont give direct answers. it doesnt help the person learn

OpenStudy (anonymous):

I get what cz1 is explaining i need to double check my answer Thank you. this all makes much more sense to me.